OpenStudy (anonymous):

5x+10y=2 solve for y. Must be written in form Y=mx+b

OpenStudy (lanhikari22):

10 y = 2 - 5x y = (2 - 5x) / 10

OpenStudy (anonymous):

To make this y=mx+b, you do this: y=(2-5x)/10 y=0.2-0.5x y=-0.5x+0.2